undeadknight26 (undeadknight26):

Choose the equation of the line passing through the point (-1, 7) and perpendicular to y = -negative one fourthx - 5. y = -4x - 11 y = -4x - 3 y = 4x + 11 y = 4x + 3

OpenStudy (shamil98):

Perpendicular lines have inverse slopes. what is the inverse of negative one over four? (-1/4)

undeadknight26 (undeadknight26):

1/-4?

OpenStudy (shamil98):

no. the inverse is the complete opposite for example. the inverse of +3 is -(1/3) so the inverse of -(1/4) is?

undeadknight26 (undeadknight26):

+(1/4)?

OpenStudy (shamil98):

no.

OpenStudy (shamil98):

The inverse of 3 is -1/3 so based of this the inverse of -1/4 is??

OpenStudy (shamil98):

the inverse of -1/3 is also 3. btw.

undeadknight26 (undeadknight26):

4 oh i get it...

OpenStudy (shamil98):

Yes.

OpenStudy (shamil98):

Now plug in those coordinates (-1,7) and your slope of 4 into y=mx + b and solve for b.

OpenStudy (shamil98):

7 = 4(-1) + b

undeadknight26 (undeadknight26):

+1 to both sides 7 = 4(-1) + b +1 +1 8 = 4 + b -4 -4 4 = b?

OpenStudy (shamil98):

No.

OpenStudy (shamil98):

4(-1) = 4 x -1 ..

OpenStudy (shamil98):

7 = -4 + b does that help?

undeadknight26 (undeadknight26):

yes... so b = 11?

OpenStudy (shamil98):

yes. know put everything together you have b = 11, m = 4 y = mx + b

undeadknight26 (undeadknight26):

y = 4x +11 so its C?

OpenStudy (shamil98):

yep